Consider a particle moving in a central force field


           f(u) = - ku2


using standard notation u =1/r. The quantity 


d2u/dθ2 + u


(a) is proportional to u2


(b) is proportional to u


(c) is a constant


(d) is proportional to 1/u


(e) none of these


1
Expert's answer
2021-03-08T08:18:13-0500
"\frac{d^2u}{d\theta^2}+u=-\frac{f(u)}{mh^2u^2}\\
\frac{d^2u}{d\theta^2}+u=-\frac{ - ku^2}{mh^2u^2}
\\
\frac{d^2u}{d\theta^2}+u=\frac{ k}{mh^2}=const"

Answer: (c) is a constant